OpenStudy (anonymous):

Solve the equations: 5x^2+4y^2=16 -----eqn 1 13x^2-5y^2=57-----eqn 2

OpenStudy (anonymous):

anyone there

OpenStudy (amistre64):

prolly try elimination, or a cramer rule

OpenStudy (amistre64):

the y parts are prolly easiest to get rid of visually; multiply the top by 5 and the bottom by 4, than add em up

OpenStudy (anonymous):

first i multiplied eqn 1 by 5 and 2 by 4 to get

OpenStudy (anonymous):

and then added both equations

OpenStudy (anonymous):

and the result is 77x=308

OpenStudy (anonymous):

then the value of x will be coming+2,-2

OpenStudy (amistre64):

that seems reasonable :) 5(5)x^2+4(5)y^2=16(5) 13(4)x^2-5(4)y^2=57(4) ------------------------ x^2(25+52) = 80+228 it looks like youve already solved it for the most part

OpenStudy (anonymous):

then i substituted the value of x in eqn 1

OpenStudy (anonymous):

i dont know the last part so that is why i am giving you all hints to solve last part of the question

OpenStudy (anonymous):

are you allowed complex solutions?

OpenStudy (amistre64):

5x^2 + 4y^2 = 16 the last part is to finish it up ....

OpenStudy (anonymous):

on substituting in eqn 1 we will get the value of y i.e, what??????

OpenStudy (amistre64):

if you did all that work by yourself above the rest should be a walk in the park for you 5(2)^2 + 4y^2 = 16 5(4) + 4y^2 = 16 20 + 4y^2 = 16 4y^2 = -4 y^2 = -1

OpenStudy (anonymous):

first equation is an ellipse second is a hyperbola they do not intersect

OpenStudy (anonymous):

since as amistre wrote you get \(y^2=-1\) which is certainly not possible with real numbers

OpenStudy (anonymous):

it will be -1=i am i right

OpenStudy (anonymous):

iota

OpenStudy (amistre64):

if you are allowed numbers that fall outside of the set of "Real", then yes, it would be "i" otherwise there are no "Real" solutions

OpenStudy (anonymous):

it is allowed

OpenStudy (anonymous):

it will be +i and -i on squaring on both sides

OpenStudy (anonymous):

its not done yet

OpenStudy (anonymous):

now for x=-2 again the answer will be coming +i and -i

OpenStudy (amistre64):

of course it is sine (+-2)^2 = 4

OpenStudy (amistre64):

since

OpenStudy (anonymous):

thank u all i was making a silly mistake

OpenStudy (amistre64):

we all make the silly mistakes :)

OpenStudy (anonymous):

thanks its done the solution set is {(2,+i,-i),(-2,+i,-i)} thanks the book shows the answer is absolutely right